OpenStudy (anonymous):

For the function f(x) = 3x+2, find x such that f(x) = 14.

OpenStudy (anonymous):

@iGreen

hero (hero):

f(x) = 3x + 12 f(x) = 14 Therefore 14 = 3x + 12 Solve for x

OpenStudy (igreen):

Can you finish that @Shadowgirl15 ?

OpenStudy (igreen):

\(12 + 3x = 14\) Subtract the 12: \(3x = 14 - 12\)

OpenStudy (anonymous):

ill try

OpenStudy (igreen):

14 - 12 = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (anonymous):

@iGreen it the answer 2?